OpenStudy (anonymous):

What property would I use to rewrite this expression? I'm not sure let me type it below

OpenStudy (anonymous):

\[\log \sqrt[6]{25x^2} = \frac{ 1 }{ 3 } \log 5x\]

OpenStudy (anonymous):

I think it may be commutative?

OpenStudy (anonymous):

this is for log expansion rewrite your radicals as exponents

OpenStudy (anonymous):

No I need to choose a property to rewrite that expression these are my options Product Property Quotient Property Power Property Commutative Property

OpenStudy (anonymous):

most likely a power

OpenStudy (anonymous):

see how on the right-hand side your exponents are gone and has been rewritten

OpenStudy (anonymous):

ignoring the log portion \(\large (25x^2)^{\frac{1}{6}} \)

OpenStudy (anonymous):

then you follow the laws of exponents

OpenStudy (anonymous):

ok so then def power property

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

You're welcome! I hope you learned something today.

OpenStudy (anonymous):

Yeah i'll look over those links for sure